OpenStudy (anonymous):

Solve the word problem. A group of roommates decide to split the cost of a television for their house. If a is the amount spent on the television and n is the number of roommates, the formula C = a/n determines each person’s share of the cost. Each person’s share was $179.60 and 5 roommates live in the house. What was the cost of the television? A. $184.60 B. $679.60 C. $898.00 D. $900.00

OpenStudy (drummergirl3):

Ahh okay...so then we have 179.60=a5 a=179.60×5 a=898

OpenStudy (drummergirl3):

got it?? @OpenStudier0

OpenStudy (anonymous):

Not really

OpenStudy (drummergirl3):

well..

OpenStudy (freckles):

@OpenStudier0 do you know you are given that the each person should spend C=a/n where a how much the tv is and n is the number of people ? And you are asked to find a when C is 179.6 and n is 5?

OpenStudy (freckles):

you plug these in C=a/n 179.6=a/5 solve for a

OpenStudy (anonymous):

I think it would be D, not sure. Just round it up. If everyone spent around let's say 200. 200x5=1000 So rounding it up 1000 is closest to 900. D would be your answer.